OpenStudy (anonymous):

What is the slope of the line passing through the points (–3, 4) and (2, –1)? A. 1 B. –1 C. 3/5 D. -5/3

Mehek (mehek14):

\(\tt{\dfrac{-1-4}{2-(-3)}=\dfrac{-5}{5}=-1}\)

OpenStudy (anonymous):

can you help me with more?